ATC51飞usiness Applications case_ Using ABC to improve product costing Hydraulic Hoses, Inc. produces specialized industrial hoses for applications such as high-pressure hydraulics and the transference of highly corrosive materials. The company recently implemented an ABC system for three of its products and is interested in evaluating its effectiveness before converting to an ABC system for all of its products. To perform this evaluation the company has compiled data for the three products using both the traditional system and the new ABC system. The traditional 250 Chapter 5 system uses a single driver (direct materials costs). The ABC system uses a variety of cost drivers related to the activities used to produce the metal products. The three products involved in the trial run of the ABC system were D-13, K-17, and R-23. The following data relate to these products. Total Costs Allocated: Traditional Costing Selling Price per Unit Units Produced Cost per Unit: Traditional Costing Total Cost Allocated ABC Cost per Unit ABC Product $14.65 15.60 18.50 250,000 $2,100,000 140,000 ,280,000 214,500 $3,594,500 $2,000,000 8.00 8.82 $8.40 D-13 K-17 1,235,000 359,500 $3,594,500 9.14 20,000 10.73 17.98 Totals Required a. Determine the gross profit margin for each product produced based on the ABC data [(Selling b. Determine the gross profit margin for each product produced based on the traditional costing data c. Provide a possible explanation as to why the cost of R-23 increased under the ABC system while d. Suggest what action management might take with respect to the discoveries resulting from the price-ABC cost per unit) × Unit produced Selling price-Traditional cost per unit) x Unit produced] the cost D-13 decreased. ABC versus traditional costing analysis. Assume that Hydraulic Hoses expects to produce a gross profit margin on each product of at least 40 percent of the selling price

Determine the gross profit margin for each product produced based on the ABC data [(selling price - ABC cost per unit) x Units produced]

b.) Determine the gross profit margin for each product produced based on the traditional costing data [(selling price - traditional cost per unit) x units produced}

c,) Provide an explanation as to why the cost of OS-367 may have increased under the ABC system while the cost of GS-157 decreased.

d.) Suggest what action management might take with respect to the discovering resulting from the ABC versus traditional costing analysis. Assume that Drilling Innovations expects to produce a gross profit margin on each product of at least 30 percent of the selling price.
Answer #1

a.Gross Profit Margin based on ABC data:

D-13 = (14.65-8)*250,000= $1,662,500

K-17 = (15.60-8.82)*140,000= $949,200

R-23 = (18.50-17.98)*20,000 = $10,400

b.Gross Profit Margin based on Traditional Costing Data:

D-13 = (14.65-8.40)*250,000= $1,562,500

K-17 = (15.60-9.14)*140,000= $904,400

R-23 = (18.50-10.73)*20,000 = $155,400

c.Under the traditional cost system, all the overheads are absorbed on the basis of direct material cost while different drivers are used under ABC system.

Cost of R-23 increased and D-13 decreased as R-23 must be incurring lower direct material cost while absorbing higher overheads on the basis of drivers.

d.Management might discontinue R-23 since the gross profit margin is very low for the product.